OpenStudy (melstutes):

Sales at a certain department store follow the model y = 85 - 13.254x where y is the total sales in thousands of dollars and x is the number of years after 2001. What was the first year that sales fell below $50,000? a.1999 c.2004 b.1998 d.2005

OpenStudy (melstutes):

Please help

OpenStudy (muzzack):

\(\large\cal\color{violet}W\color{violet}e\color{violet}l\color{violet}c\color{violet}o\color{violet}m\color{violet}e\ ~\color{violet}T\color{violet}o\ ~Openstudy\)

OpenStudy (muzzack):

im pretty sure its B

OpenStudy (melstutes):

explain please

OpenStudy (muzzack):

y thousand dollars = $50,000 y = 50 50 = 85-13.254x 13.254x = 85-50 x = 35/13.254 x = 2.64 x ~ 3 The first year that sales fell below 50,000 would be 2001+3 = 2004 :)